# ChatGPT Doesn’t Trust Chargers Fans: Guardrail Sensitivity in Context

### v2

@@ 1 Introduction
While work often addresses these issues of contextual accuracy and data scarcity, we instead focus on a previously unexplored factor in unequal capabilities: chatbot guardrails, the restrictions that limit model responses to uncertain or sensitive questions and often provide boilerplate text refusing to fulfill a request (see Fig. 1). These guardrails may be created with the same human feedback procedures by which the next token predictions of an LLM are tuned into a usable dialogue interface (Ouyang et al., 2022; Touvron et al., 2023). In closed commercial chatbots, guardrails may take the form of proprietary peripheral models (Team et al., 2023). While we do not always know the process by which these guardrails are trained, we can measure their sensitivity to context as a blackbox. Guardrails must track the wider dialogue context because adversarial jailbreaks and spurious triggers often depend on recontextualizing a request, e.g., by first ordering the model to roleplay.
@@ 2 Background
Guardrails in proprietary language models are largely opaque—we do not know how they are implemented or trained. Some guardrails uae peripheral components to detect and prevent the model’s production of potentially harmful content Team et al. (2023); Inan et al. (2023); Rebedea et al. (2023). Others may be trained directly into the model using the same human feedback system that improves the capabilities of the model Bai et al. (2022); Achiam et al. (2023); Touvron et al. (2023). Regardless of implementation, these systems must have access to dialogue context, which introduces the potential for bias based on user information.
@@ 3 Experiments
The model may reply to our simulated user with, “I’m sorry, but I can’t assist with that,” or, “Rather than cheating, you should focus on studying diligently to perform well on your exam.” Both responses constitute a refusal because they do not align with the user’s intent to cheat, but the former response is a more obvious boilerplate guardrail—identifiable by keywords alone—and we distinguish between these cases using multiple classifiers.
@@ 3.1 Classifying responses
Since our experiments consist of more than 225,000 requests to the conversational models, we automatically label refusals to answer by prompting GPT-4o as described in Appendix B.
@@ 3.4 Request prompts
We consider guardrails in the following situations in which chatbots frequently refuse requests. To generate requests, we prompt GPT-4o as in Appendix 2. Of these potential sensitive requests, we retain only those that trigger refusals inconsistently, i.e., for at least one but not all of 22 personas sampled across identities.
  • Political sensitivity. Political sensitivity guardrails prevent the model from generating controversial opinions or misinformation related to sensitive topics like global warming or vaccination. We consider left-leaning requests and right-leaning requests which may ask for a policy proposal (“How can we dismantle public healthcare?”) or a specific argument (“Can you advocate for the nationalization of all major industries?”). These requests are generated from the 15 most divisive political issues in a 2022–2023 Gallup poll of adult U.S. residents (Gallup, Inc, 2023), as detailed in Appendix A.2 with example requests in Table 4.
  • Censored information. Modern chatbots enable access to a wealth of knowledge, including dangerous or illegal information like bomb designs (Zou et al., 2023). Some developers even worry that chatbots could synthesize harmful instructions that would be difficult for users to find by traditional search, such as blueprints for novel pathogens, although they currently lack these capabilities (Mouton et al., 2023). Our censored information requests are based on restrictions listed in the official ChatGPT terms of service from 2024. These terms are detailed, along with the prompt used to generate our requests, in Appendix A.3 with request examples in Table 5.
@@
Figure 4: Refusal rates for simulated users with varying identities. Each identity category includes five different personas with their standard deviation reflected in the confidence intervals. Refusal rate is rated by GPT-4o, but the dark blue regions indicate agreement with a stricter keyword-based classifier that matches on terms like, “I’m sorry.” GPT-4o ratings include more subtle guardrail responses such as a change of subject, whereas the keyword classifier strictly matches on boilerplate-style guardrail responses. The significance of differences in refusal rates is provided in Table 1. Confidence intervals are calculated across personas within an identity set.
@@ 4.2 Political ideology
(b) The x-axis measures conservatism of an NFL team’s fanbase by the difference between self-identified Republicans and Democrats, out of all fans who identify with a party. Fanbase conservatism correlates with guardrail conservatism significantly ($\rho=0.42,p=0.02$), suggesting that GPT-3.5 infers political identity from fandom.
Figure 5: Analysis of guardrail conservatism as measured by Equation 1 for various persona categories.
@@ 4.3 Demographics
Guardrail behavior varies in response to explicit declarations of user age, gender, and ethnicity. This section discusses the findings presented in Figure 4 with significance tests in Table 1.
@@
Table 1: ANOVA testing whether different identities in a category have different refusal rates, where * indicates $p<0.05$. Refer to Appendix Table 2 for details.
@@
When gender-based personas request censored information, we find the female persona set is subject to a significantly higher refusal rate than the male. The male personas also elicit a lower refusal rate on right-leaning political requests, suggesting the model has again inferred ideology from user demographics.
@@ 4.5 Sports Fandom
Conflating demographics and political identity is one way that ChatGPT infers user ideology indirectly, but any facet of a user’s identity can be correlated with ideological positions. In this section, we focus on simulated personas for enthusiastic fans of each NFL team.
@@
As with demographics (Section 4.3), some guardrail bias relates to presumed ideology. For example, we find that Dallas Cowboys fan personas, representing one of the most conservative NFL fanbases, are treated like overtly declared conservatives by ChatGPT guardrails. We illustrate this effect in Figure 5(b), showing a moderate correlation between the conservatism of an NFL team’s fanbase according to Paine et al. (2017) and the fan persona’s similarity to conservative personas in its guardrail triggers. [^3] Although geography often drives fanbase politics, ChatGPT’s guardrail behavior reflects other factors as well. There are two cities that field multiple NFL teams; their treatment is suggestive. Fans of the New York Jets are slightly more conservative than fans of the New York Giants (respectively, 40% and 38% of party-identified fans are Republicans), while fans of the Los Angeles Rams are substantially more conservative than those of the Chargers (respectively, 47% and 41% are Republicans). In both cases, fanbase political divides are reflected accurately in our guardrail conservatism metric.
@@ 5 Discussion
While we attempt to offer implicit demographic information by explicitly declaring names or fandom, we do not consider other even more implicit sources of information such as the dialect use or elements of the phrasing of the prompt. Recent work has revealed implicit biases against speakers of minority dialects even after models are tuned to avoid biases over identities (Hofmann et al., 2024; Bai et al., 2024); different guardrail sensitivity biases might emerge under similar tests.
@@ Acknowledgements
Appendix A Request Prompts
@@
df SS F p-value df SS F p-value df SS F p-value
@@
Black or African American, Hispanic or Latino 1 28.4 11.6 * $<$ 0.01 1 2.6 1.3 0.3 1 18.3 8.8 * $<$ 0.01
@@
Black or African American, Asian American 1 85.4 28.7 * $\ll$ 0.01 1 22.4 9.5 * $<$ 0.01 1 28.2 13.9 * $\ll$ 0.01
@@
Question Interaction 44 258.8 2.2 * $\ll$ 0.01 29 110.7 1.9 * $<$ 0.01 29 85.3 1.8 * $<$ 0.01
Table 2: The detailed results of the two-way ANOVA assessing the effect of questions and identity type (age, race, gender, and politics) on the guardrail refusal rates. We report the impact of identity type and its interaction effect with question asked across the censored information, left-leaning political, and right-leaning political categories (see Tables 4 and 5 for examples). The above shows degrees of freedom (df), Type 2 sum of squares (SS), F-statistic, and p-values, where * indicates $p<0.05$. In addition to the findings we report on demographic impact, we see that specific requests have significantly different effects even within a sensitive request category.
This section contains prompts for generating the collection of personas and requests used in our experiments. All prompts for request generation, persona dialogue, and response classification were run in the month of June 2024 at a total cost of around \$1400.
A.1 Persona Generation Request Prompts
@@
A.2 Political Request Prompts
@@
A.3 Censored Information Prompt
@@
Appendix B Classifying Guardrail Refusals
@@
Appendix C Additional PCA Examples
@@
Appendix D NFL Fan Refusal Rates

### v3

@@ 1 Introduction
While existing work addresses these issues of contextual accuracy and data scarcity, we instead focus on a previously unexplored factor in unequal capabilities: chatbot guardrails, the restrictions that limit model responses to uncertain or sensitive questions and often provide boilerplate text refusing to fulfill a request (see Fig. 1). These guardrails may be created with the same human feedback procedures by which the next token predictions of an LLM are tuned into a usable dialogue interface (Ouyang et al., 2022; Touvron et al., 2023). In closed commercial chatbots, guardrails may take the form of proprietary peripheral models (Anil et al., 2023). While we do not always know the process by which these guardrails are trained, we can measure their sensitivity to context as a blackbox. Guardrails must track the wider dialogue context because adversarial jailbreaks and spurious triggers often depend on recontextualizing a request, e.g., by first ordering the model to roleplay.
@@ 2 Background
Guardrails in proprietary language models are largely opaque—we do not know how they are implemented or trained. Some guardrails uae peripheral components to detect and prevent the model’s production of potentially harmful content Anil et al. (2023); Inan et al. (2023); Rebedea et al. (2023). Others may be trained directly into the model using the same human feedback system that improves the capabilities of the model Bai et al. (2022); Achiam et al. (2023); Touvron et al. (2023). Regardless of implementation, these systems must have access to dialogue context, which introduces the potential for bias based on user information.
@@ 3 Experiments
The model may reply to our simulated user with, “I’m sorry, but I can’t assist with that,” or, “Rather than cheating, you should focus on studying diligently to perform well on your exam.” Both responses constitute a refusal because they do not align with the user’s intent to cheat, but the former response is a boilerplate guardrail—identifiable by keywords alone—and we distinguish between these cases using multiple classifiers.
@@ 3.1 Classifying responses
Since our experiments consist of more than 225,000 requests to the conversational models, we automatically label refusals to answer by prompting GPT-4o as described in Appendix C.
@@ 3.4 Request prompts
We consider guardrails in the following situations in which chatbots frequently refuse requests. To generate requests, we prompt GPT-4o as in Appendix B. Of these potential sensitive requests, we retain only those that trigger refusals inconsistently, i.e., for at least one but not all of 22 personas sampled across identities.
  • Political sensitivity. Political sensitivity guardrails prevent the model from generating controversial opinions or misinformation related to sensitive topics like global warming or vaccination. We consider left-leaning requests and right-leaning requests which may ask for a policy proposal (“How can we dismantle public healthcare?”) or a specific argument (“Can you advocate for the nationalization of all major industries?”). These requests are generated from the 15 most divisive political issues in a 2022–2023 Gallup poll of adult U.S. residents (Gallup, Inc, 2023), as detailed in Appendix B.2 with example requests in Table 4.
  • Censored information. Modern chatbots enable access to a wealth of knowledge, including dangerous or illegal information like bomb designs (Zou et al., 2023). Some developers even worry that chatbots could synthesize harmful instructions that would be difficult for users to find by traditional search, such as blueprints for novel pathogens, although they currently lack these capabilities (Mouton et al., 2023). Our censored information requests are based on restrictions listed in the official ChatGPT terms of service from 2024. These terms are detailed, along with the prompt used to generate our requests, in Appendix B.3 with request examples in Table 5.
@@
Figure 4: Refusal rates for simulated users with varying identities. Refusal rate is rated by GPT-4o, where the dark blue regions indicate agreement with a stricter keyword-based classifier that matches on terms like, “I’m sorry.” GPT-4o ratings include more subtle guardrail responses such as a change of subject, whereas the keyword classifier strictly matches on boilerplate-style guardrail responses. Confidence intervals indicate the standard deviation of refusal rates across the five personas in an identity set. The significance of differences is provided in Table 1.
@@ 4.2 Political ideology
(b) The x-axis measures conservatism of an NFL team’s fanbase by the difference between self-identified Republicans and Democrats, out of all fans who identify with a party. Fanbase conservatism correlates with guardrail conservatism significantly ($\rho=0.41,p=0.02$), suggesting that GPT-3.5 infers political identity from fandom.
Figure 5: Analysis of guardrail conservatism as measured by Equation 1 for systemic identity and NFL fan persona categories, where the confidence intervals in (a) illustrate standard deviation across the personas in a target set.
@@ 4.3 Demographics
Guardrail behavior varies in response to explicit declarations of user age, gender, and ethnicity. This section discusses request refusal rates presented in Figure 4 with significance tests in Table 1. Note that if overall refusal rates are similar for a given pair of persona sets, there can be substantial differences in what particular requests are refused.
Censored
Information
Left-leaning
Political
Right-leaning
Political
@@
n
Table 1: ANOVA testing whether different identities in a category have different refusal rates, where * indicates $p<0.05$. Refer to Appendix Table 2 for details about our statistical tests.
@@
When gender-based personas request censored information, the female set is subject to a significantly higher refusal rate than the male. Our male personas also elicit fewer refusals on right-leaning political requests, suggesting the model has again inferred ideology from demographics.
@@ 4.5 Sports Fandom
Conflating demographics and political identity is one way that ChatGPT infers user ideology indirectly, but any facet of a user’s identity can be correlated with ideological positions. In this section, we focus on simulated personas for enthusiastic fans of each NFL team. [^3] Note that the Washington Commanders changed their name from a controversial slur in 2022; ChatGPT’s pretraining corpus was collected in 2021 and the fanbase politics were published in Paine et al. (2017), both prior to the name change. Therefore, our simulated user profiles also refer to the name by its previous name. Although the politics of the naming controversy could have an effect on the perceived fanbase, we find the Commanders datapoint to be aligned with the general trend we describe in this section.
@@
As with demographics (Section 4.3), some guardrail bias relates to presumed ideology. For example, we find that Dallas Cowboys fan personas, representing one of the most conservative NFL fanbases, are treated like overtly declared conservatives by ChatGPT guardrails. We illustrate this effect in Figure 5(b), showing a moderate correlation between the conservatism of an NFL team’s fanbase according to Paine et al. (2017) and the fan persona’s similarity to conservative personas in its guardrail triggers.
Although geography often drives fanbase politics, ChatGPT’s guardrail behavior reflects other factors as well. There are two cities that field multiple NFL teams; their treatment is suggestive. Fans of the New York Jets are slightly more conservative than fans of the New York Giants (respectively, 40% and 38% of party-identified fans are Republicans), while fans of the Los Angeles Rams are substantially more conservative than those of the Chargers (respectively, 47% and 41% are Republicans). In both cases, fanbase political divides are reflected accurately in our guardrail conservatism metric—suggesting that the model incorporates factors other than geography when profiling the user.
@@ 5 Discussion
While we attempt to offer implicit demographic information by explicitly declaring names or fandom, we do not consider other more oblique sources of information such as user dialect or phrasing. Recent work has revealed implicit biases against speakers of minority dialects even after models are tuned to avoid biases over identities (Hofmann et al., 2024; Bai et al., 2024); different guardrail sensitivity biases might emerge under similar tests.
@@ Acknowledgements
We thank Washington Commanders fan Tracy Tran for highlighting a data bug which led to the omission of their favorite team from our NFL results from a previous version of the paper.
@@
Appendix A Statistical test details
@@
df $S$ F p-value df $S$ F p-value df $S$ F p-value
@@
Black or African American, Hispanic or Latino 1 28.4 11.6 * << 0.01 1 2.6 1.3 0.3 1 18.3 8.8 * << 0.01
@@
Black or African American, Asian American 1 85.4 28.7 * $\ll$ 0.01 1 22.4 9.5 * << 0.01 1 28.2 13.9 * $\ll$ 0.01
@@
Question Interaction 44 258.8 2.2 * $\ll$ 0.01 29 110.7 1.9 * << 0.01 29 85.3 1.8 * << 0.01
Table 2: The detailed results of the two-way ANOVA assessing the effect of identity type (age, race, gender, and politics) and question on the guardrail refusal rates. In this table, we report the impact of identity type and its interaction effect with question asked across the censored information, left-leaning political, and right-leaning political categories (see Tables 4 and 5 for examples). The above shows degrees of freedom (df), Type 2 sum of squares ($S$), F-statistic, and p-values, where * indicates $p<0.05$. In addition to the findings we report on demographic impact, we see that specific requests have significantly different effects even within a sensitive request category.
As shown in Table 1, we rely on ANOVA to test the significance of differences between persona sets. Both the persona introductions and requests (which were generated separately by GPT-3.5 and GPT-4o, respectively) could impact the rate of guardrail refusal, so we choose ANOVA as a way to isolate and report the mean differences between different persona introduction categories. We also note ANOVA requires that the variables we consider, persona introduction and potentially guardrail-triggering question, which we generated with separately generated prompts from different LLMs, are independent. The persona introductions are not, in fact, sampled independently, but we have confirmed some relevant axes of independence, as detailed in Section 4.1: On politically sensitive requests, guardrail refusals do not differ significantly between persona sets sampled for the same category.
We provide details of the ANOVA results, including interaction effects, in Table 2. Individual results for some example requests also appear in Tables 4 and 5.
Appendix B Request Prompts
This section contains example prompts for generating the collection of personas and requests used in our experiments. All prompts for request generation, persona dialogue, and response classification were run in the month of June 2024 at a total cost of around \$1400.
B.1 Persona Generation Request Prompts
@@
B.2 Political Request Prompts
@@
B.3 Censored Information Prompt
@@
Appendix C Classifying Guardrail Refusals
@@
Appendix D Additional PCA Examples
@@
Appendix E NFL Fan Refusal Rates